OpenStudy (goformit100):

Show that the sum of the vectors represented by the sides vector AB, vector DC of a quadrilateral ABCD is equivalent to the sum of the vectors represented by the diagonals vector AC and vector DB.

OpenStudy (goformit100):

@satellite73

OpenStudy (goformit100):

@Callisto @RadEn @Reaper534 @RoRolando- @Reallywanttolearn

OpenStudy (goformit100):

Provide me a diagram

OpenStudy (goformit100):

@audrielynn

OpenStudy (goformit100):

Help

OpenStudy (anonymous):

 now AC+DB=(AO+OC)+(DO+OB) =(AO+OB)+(DO+OC) =AB+DC PROVED